OpenStudy (anonymous):

A ball is thrown vertically upward at an initial velocity of 32fps. The distance s (in feet) of the ball from the ground after t seconds is s=32t-16t^2. A) At what time t will the ball strike the ground? B) For what time t is the ball more than 12 feet above the ground?

OpenStudy (cwrw238):

when ball strikes the ground s = 0 0 = 32t - 16t^2 one value of t will be 0 which corresponds to the time the ball was thrown up

OpenStudy (cwrw238):

the other value of t is what you want

OpenStudy (anonymous):

t=8?

OpenStudy (cwrw238):

16t^2 - 32t = 0 16t(t - 2) = 0 t = 0 or t = 2

OpenStudy (cwrw238):

2 seconds

OpenStudy (anonymous):

so if im going to figure out when its above 12 ft how would I plug it into my equation?

OpenStudy (cwrw238):

i'm looking at this again - 2 seconds seem a very short time

OpenStudy (cwrw238):

- no its is correct

OpenStudy (cwrw238):

well you can first work out the time taken to get to 12 feet from 12 = 32t - 16t^2 16t^2- 32t + 12 = 0 4t^2 - 8t + 3 = 0 t = 0.5 ,1.5 0.5 secs as it climbs then 1.5 secs on the way down

OpenStudy (cwrw238):